# Algorithmic contiguity from low-degree conjecture and applications in correlated random graphs

### v1

@@ Abstract
In this paper, assuming a natural strengthening of the low-degree conjecture, we provide evidence of computational hardness for two problems: (1) the (partial) matching recovery problem in the sparse correlated Erdős-Rényi graphs $\mathcal{G}(n,q;\rho)$ when the edge-density $q=n^{-1+o(1)}$ and the correlation $\rho<\sqrt{\alpha}$ lies below the Otter’s threshold, solving a remaining problem in [15]; (2) the detection problem between the correlated sparse stochastic block model $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$ and a pair of independent stochastic block models $\mathcal{S}(n,\tfrac{\lambda s}{n};k,\epsilon)$ when $\epsilon^{2}\lambda s<1$ lies below the Kesten-Stigum (KS) threshold and $s<\sqrt{\alpha}$ lies below the Otter’s threshold, solving a remaining problem in [9].
@@ 1 Introduction
Graph matching, also referred to as network alignment, is the problem of identifying a bijection between the vertex sets of two graphs that maximizes the number of common edges. When the two graphs are exactly isomorphic to each other, this problem reduces to the classical graph isomorphism problem, for which the best known algorithm runs in quasi-polynomial time [1]. In general, graph matching is an instance of the quadratic assignment problem [7], which is known to be NP-hard to solve or even approximate [38].
Motivated by real-world applications (such as social network de-anonymization [45] and computational biology [51]) as well as the need to understand the average-case computational complexity, recent research has focused on developing theoretical foundations and efficient algorithms for graph matching under statistical models. These models assume that the two graphs are randomly generated with correlated edges under a hidden vertex correspondence, and a canonical model among them is the following correlated random graph model. For any integer $n$, denote by $\operatorname{U}=\operatorname{U}_{n}$ the set of unordered pairs $(i,j)$ with $1\leq i\neq j\leq n$.
@@
Given an integer $n\geq 1$, for $(i,j)\in\operatorname{U}_{n}$ let $J_{i,j}$ and $K_{i,j}$ be independent Bernoulli variables with parameter $s$. In addition, let $\pi_{*}$ be an independent uniform permutation on $[n]=\{1,\dots,n\}$. Then, we define a triple of correlated random graphs $(G,A,B)$ as follows: first we generate $G$ independently with $\{I_{i,j},K_{i,j}\}$ and $\pi_{*}$ from a specific probability distribution over all graphs on $[n]$, and then (conditioned on $G$) we define for each $(i,j)\in\operatorname{U}_{n}$ that (note that we identify a graph with its adjacency matrix)
@@
Of particular interest in our paper are the following two special cases, namely the correlated Erdős-Rényi models and the correlated stochastic block models (SBMs).
@@
Given an integer $n\geq 1$ and two parameters $p,s\in(0,1)$, we generate a triple of correlated random graphs $(G,A,B)$ such that we first generate $G$ according to an Erdős-Rényi graph distribution $\mathcal{G}(n,p)$ (i.e., for each $(i,j)\in\operatorname{U}_{n}$ we connect $(i,j)$ in $G$ independently with probability $p$), and then generate $(A,B)$ from $G$ according to Definition 1.1. For ease of presentation, we shall reparameterize such that $q=ps$ and $\rho=\frac{s(1-p)}{1-ps}$ respectively. We will denote the marginal law of $(A,B)$ as $\mathcal{G}(n,q;\rho)$.
@@
Definition 1.4 (Correlated stochastic block models).
@@
Given an integer $n\geq 1$ and four parameters $k\in\mathbb{N},\lambda>0,\epsilon,s\in(0,1)$, for $(i,j)\in\operatorname{U}_{n}$ let $J_{i,j}$ and $K_{i,j}$ be independent Bernoulli variables with parameter $s$.We define a triple of correlated random graphs $(G,A,B)$ as follows: we first sample $G$ according to the law of a stochastic block model $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon)$ and then generate $(A,B)$ from $G$ according to Definition 1.1. We will denote the marginal law of $(A,B)$ as $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$.
@@
Two fundamental problems in the study of correlated random graph models are as follows: (1) the detection problem, which involves determining whether a given pair of graphs $(A,B)$ is sampled from a pair of correlated random graphs or from a pair of independent random graphs; (2) the matching problem, which focuses on recovering the latent matching $\pi_{*}$ from a sample $(A,B)$ from the distribution of correlated random graphs. In recent years, significant progress has been made in understanding these problems for both the correlated Erdős-Rényi model and correlated stochastic block models (SBMs). Through the collective efforts of the community, the information-theoretic thresholds for detection and matching have been fully characterized for correlated Erdős-Rényi model and partially characterized for correlated SBMs [10, 29, 54, 55, 25, 13, 14, 47, 28]. Additionally, various efficient detecting and matching algorithms have been developed with performance guarantees [11, 4, 18, 20, 21, 24, 26, 27, 39, 40, 42, 41, 16, 17, 8, 9]. Notably, like many other inference tasks in high-dimensional statistics [57, 48, 37, 22], these problems appear to exhibit information-computation gaps. Specifically, for certain ranges of the correlation strength, detection/matching is information theoretically possible but no efficient algorithm is known to achieve these tasks. We now focus on the algorithmic side of these problems as they are more relevant to our work. Indeed, it has been shown that many inference tasks in the correlated random graph models exhibits sharp algorithmic phase transitions, as we elaborate below:
@@
  • For the detection problem between a pair of correlated Erdős-Rényi models $\mathcal{G}(n,q;\rho)$ and two independent Erdős-Rényi models $\mathcal{G}(n,q)$, we focus on the sparse regime where $q=n^{-1+o(1)}$. In this regime, on the one hand, it was shown in [42] that when $\rho>\sqrt{\alpha}$ where $\alpha\approx 0.338$ is the Otter’s constant [46], there is an efficient algorithm that strongly distinguish these two models; on the other hand, it was shown in [15] that when $s<\sqrt{\alpha}$ there are evidences suggesting that all algorithms based on low-degree polynomials fail to strongly distinguish these two models.
@@
  • For the detection problem between a pair of correlated SBMs $\mathcal{S}(n,\frac{\lambda}{n};k,\epsilon;s)$ and two independent Erdős-Rényi graphs $\mathcal{G}(n,\frac{\lambda s}{n})$, we focus on the constant degree regime where $\lambda=O(1)$. In this regime, on the one hand, it was shown in [9] that when $s>\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ where $\alpha\approx 0.338$ is the Otter’s constant and $\tfrac{1}{\epsilon^{2}\lambda}$ is the Kesten-Stigum threshold [33], there is an efficient algorithm that strongly distinguish these two models; on the other hand, it was also shown in [9] that when $s<\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ there are evidences suggesting that all algorithms based on low-degree polynomials fail to strongly distinguish these two models.
@@
The lower bound in the aforementioned results explored inherent computational barriers from the perspective of the low-degree polynomial framework. Indeed, it has been proved that the class of low-degree polynomial algorithms is a useful proxy for computationally efficient algorithms, in the sense that the best-known polynomial-time algorithms for a wide variety of high-dimensional inference problems are captured by the low-degree class; see e.g. [31, 30, 50, 37]. However, these aforementioned results suffer from two significant limitations, which we now discuss. Firstly, the aforementioned result cannot provide the evidence that partial matching recovery (that is, recover a positive fraction of the coordinates of the latent matching $\pi_{*}$) is impossible by efficient algorithms although efficient detection is already ruled out. This limitation arises because there is no evident reduction from partial matching to detection. Secondly, (in the case of correlated SBMs) they are only able to establish the computation threshold on the detection problem between the correlated model and a pair of independent Erdős-Rényi graphs. These limitations motivate two natural questions:
@@
    Can we provide the evidences that partial matching recovery is impossible (by efficient algorithms) in the same parameter regime where detection is impossible (by efficient algorithms)?
@@
The aim of this paper is to find (partial) answers for these two problems. Our main result can be informally summarized as follows:
@@
Assuming a natural strengthening of the low-degree conjecture (see Conjecture 2.7 for its precise meaning), we have
@@
    For the correlated Erdős-Rényi model $\mathcal{G}(n,q,\rho)$, when $q=n^{-1+o(1)}$ and $\rho<\sqrt{\alpha}$ it is impossible to recover a positive fraction of the coordinates of $\pi_{*}$ with positive probability by efficient algorithms.
@@
    For the correlated stochastic block models $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$, when $\lambda=O(1)$ and $s<\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ it is impossible to recover a positive fraction of the coordinates of $\pi_{*}$ with probability tending to $1$ as $n\to\infty$ by efficient algorithms.
@@
    For the correlated stochastic block models $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$, when $s<\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ it is impossible to strongly distinguish this model and a pair of independent SBMs $\mathcal{S}(n,\tfrac{\lambda s}{n};k,\epsilon)$, provided that the average degree $\lambda s$ is sufficiently large.
@@
Note that when $\rho>\sqrt{\alpha}$ (respectively, $s>\sqrt{\alpha}$), the results in [41, 24, 26, 27] shows (or naturally extends to show) that there exists an efficient algorithm that achieves partial recovery of $\pi_{*}$ in a pair of correlated Erdős-Rényi models (respectively, correlated SBMs) with probability tending to $1$ as $n\to\infty$. Thus, Items (1) and (2) in Theorem 1.6 is tight in some sense and the algorithmic partial recovery threshold is indeed given by $\sqrt{\alpha}$.
@@
From Item (3) in Theorem 1.6 we see that for the correlated SBMs $(A,B)$, when marginally both $A$ and $B$ are below the KS-threshold, there is no efficient algorithm that strongly distinguish $(A,B)$ from a pair of independent stochastic block models when the correlation $s<\sqrt{\alpha}$. Since the result in [42] extends naturally to the case of stochastic block models which provides an efficient algorithm that strongly distinguish these two models when $s>\sqrt{\alpha}$, we see that in the subcritical regime (i.e., when marginally both $A$ and $B$ are below the KS-threshold) the algorithmic correlation detection threshold is given by $\sqrt{\alpha}$. On the contrary, in the supercritical regime where $A$ and $B$ are above the KS-threshold, we believe that the correlation detection threshold should be strict lower than $\sqrt{\alpha}$.
@@ 1.1 Key challenges and innovations
In this subsection, we first briefly discuss our approach of showing Theorem 1.6 and some conceptual innovations behind it. Our idea can be summarized as follows:
@@
    We will show that the low-degree advantage between $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ and $\mathbb{Q}$ is bounded by an absolute constant. Then, from the standard low-degree conjecture (see Conjecture 2.2 for details) these two measures cannot be strongly distinguished by efficient algorithms. Thus, since $g_{i}$ is “not small” under the measure $\mathbb{P}(\cdot\mid\pi_{*}(i)=1)$ (as it should approximate $\mathbf{1}_{\{\pi_{*}(1)=i\}}$ in some sense) we expect that $g_{i}$ should also be “not small” under $\mathbb{Q}$.
@@
    Define $g=g_{1}+\ldots+g_{n}$. We then have a statistics $g$ that can be efficiently calculated, and (I) under $\mathbb{Q}$ we expect that $g$ is “large” as it is the sum of $n$ “not small” terms; (II) under $\mathbb{P}$ we expect that $g$ is “small” as each $g_{i}$ should approximate $\mathbf{1}_{\{\pi_{*}(1)=i\}}$ in some sense. Thus, the statistics $g$ accumulates more signals than all low-degree polynomials, which violates the low-degree conjecture.
@@
    As for Item (3), denote $\mathbb{P}$ to be the law of correlated SBMs and $\mathbb{Q}$ of independent SBMs. Also denote $\widetilde{\mathbb{Q}}$ to be the law of independent Erdős-Rényi graphs. As it was already shown in [9] that $\mathbb{P}$ and $\mathbb{Q}$ cannot be strongly distinguished, (non-rigorously speaking) we only need to show that $\mathbb{Q}$ and $\widetilde{\mathbb{Q}}$ are also “indistinguishable” in some sense.
However, there are certain obstacles when implementing the above ideas, as we shall discuss below. In Step (2) we need to “transfer” the behavior of a statistics $g_{i}$ under $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ to its behavior under $\mathbb{Q}$. If the low-degree advantage between $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ and $\mathbb{Q}$ is given by $1+o(1)$, then the standard low-degree conjecture implies that we cannot distinguish between $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ and $\mathbb{Q}$ better than random by efficient algorithms. Consequently, we expect the behavior of $g_{i}$ to be “almost identical” under both $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ and $\mathbb{Q}$. However, in our case the low-degree advantage is just bounded by a large (but fixed) constant. This weaker condition means that the standard low-degree conjecture only rules out the possibility to distinguish these two measures with vanishing errors, leaving room for non-negligible distinctions. Similar issues also arise in Step (4), in which simply showing the low-degree advantage between $\mathbb{P}$ and $\mathbb{Q}$ and the low-degree advantage between $\mathbb{Q}$ and $\widetilde{\mathbb{Q}}$ is bounded is not strong enough for our goal.
@@
Another difficulty is that reducing between different inference tasks requires a framework that better constrains the behavior of efficiently computable estimators. To address this, we introduce a natural refinement of the low-degree conjecture: we posit that low-degree polynomials achieve signal-to-noise ratios that are at least as good as those of any efficient algorithm (see Conjecture 2.7 for a more precise description). While similar modifications to the low-degree conjecture have been explored in [44], our motivations differ significantly. Their work focused on characterizing the limits of precise error for all computationally feasible algorithms, whereas our goal is to enable reductions between distinct inference tasks.
1.2 Notation

### v2

@@ Abstract
In this paper, assuming a natural strengthening of the low-degree conjecture, we provide evidence of computational hardness for two problems: (1) the (partial) matching recovery problem in the sparse correlated Erdős-Rényi graphs $\mathcal{G}(n,q;\rho)$ when the edge-density $q=n^{-1+o(1)}$ and the correlation $\rho<\sqrt{\alpha}$ lies below the Otter’s threshold, this resolves a remaining problem in [DDL23+]; (2) the detection problem between a pair of correlated sparse stochastic block model $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$ and a pair of independent stochastic block models $\mathcal{S}(n,\tfrac{\lambda s}{n};k,\epsilon)$ when $\epsilon^{2}\lambda s<1$ lies below the Kesten-Stigum (KS) threshold and $s<\sqrt{\alpha}$ lies below the Otter’s threshold, this resolves a remaining problem in [CDGL24+].
@@ 1 Introduction
Graph matching, also referred to as network alignment, is the problem of identifying a bijection between the vertex sets of two graphs that maximizes the number of common edges. When the two graphs are exactly isomorphic to each other, this problem reduces to the classical graph isomorphism problem, for which the best known algorithm runs in quasi-polynomial time [Bab16]. In general, graph matching is an instance of the quadratic assignment problem [BCPP98], which is known to be NP-hard to solve or even approximate [MMS10]. Motivated by real-world applications (such as social network de-anonymization [NS08] and computational biology [SXB08]) as well as the need to understand the average-case computational complexity, recent research has focused on developing theoretical foundations and efficient algorithms for graph matching under statistical models. These models assume that the two graphs are randomly generated with correlated edges under a hidden vertex correspondence, and a canonical model among them is the following correlated random graph model. For any integer $n$, denote by $\operatorname{U}=\operatorname{U}_{n}$ the set of unordered pairs $(i,j)$ with $1\leq i\neq j\leq n$.
@@
Given an integer $n\geq 1$, for $(i,j)\in\operatorname{U}$ let $J_{i,j}$ and $K_{i,j}$ be independent Bernoulli variables with parameter $s$. In addition, let $\pi_{*}$ be an independent uniform permutation on $[n]=\{1,\dots,n\}$. Then, we define a triple of correlated random graphs $(G,A,B)$ as follows: first we generate $G$ independently with $\{I_{i,j},K_{i,j}\}$ and $\pi_{*}$ from a specific probability distribution over all graphs on $[n]$, and then (conditioned on $G$) we define for each $(i,j)\in\operatorname{U}$ that (note that we identify a graph with its adjacency matrix)
@@
Of particular interest in our paper are the following two special cases, namely the correlated Erdős-Rényi model and the correlated stochastic block model (SBM).
@@
Given an integer $n\geq 1$ and two parameters $p,s\in(0,1)$, we generate a triple of correlated random graphs $(G,A,B)$ such that we first generate $G$ according to an Erdős-Rényi graph distribution $\mathcal{G}(n,p)$ (i.e., for each $(i,j)\in\operatorname{U}$ we connect $(i,j)$ in $G$ independently with probability $p$), and then generate $(A,B)$ from $G$ according to Definition 1.1. For ease of presentation, we shall reparameterize such that $q=ps$ and $\rho=\frac{s(1-p)}{1-ps}$ respectively. We will denote the marginal law of $(A,B)$ as $\mathcal{G}(n,q;\rho)$.
@@
Definition 1.4 (Correlated stochastic block model).
@@
Given an integer $n\geq 1$ and four parameters $k\in\mathbb{N},\lambda>0,\epsilon,s\in(0,1)$, we define a triple of correlated random graphs $(G,A,B)$ as follows: we first sample $G$ according to the law of a stochastic block model $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon)$ and then generate $(A,B)$ from $G$ according to Definition 1.1. We will denote the marginal law of $(A,B)$ as $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$.
@@
Two fundamental problems in the study of correlated random graph model are as follows: (1) the detection problem, which involves determining whether a given pair of graphs $(A,B)$ is sampled from a pair of correlated random graphs or from a pair of independent random graphs; (2) the matching problem, which focuses on recovering the latent matching $\pi_{*}$ from a sample $(A,B)$ from the distribution of correlated random graphs. In recent years, significant progress has been made in understanding these problems for both the correlated Erdős-Rényi model and correlated stochastic block models. Through the collective efforts of the community, the information-theoretic thresholds for detection and matching have been fully characterized for correlated Erdős-Rényi model [CK17+, HM23, WXY22, WXY23, GML21, DD23a, DD23b, Du25+] and partially characterized for correlated SBMs [RS21, GRS22]. Additionally, various efficient detecting and matching algorithms have been developed with performance guarantees [DCKG19, BCL+19, DMWX21, FMWX23a, FMWX23b, GM20, GML24, GMS24, MRT21, MRT23, MWXY24, MWXY23, DL22+, DL23+, CR24, CDGL24+, Li24+]. Notably, like many other inference tasks in high-dimensional statistics [ZK16, RSS19, KWB22, Gam21], these problems appear to exhibit information-computation gaps. Specifically, for certain ranges of the correlation strength, detection or matching is information theoretically possible but no efficient algorithm is known to achieve these tasks. We now focus on the algorithmic side of these problems as they are more relevant to our work. Indeed, it has been shown that many inference tasks in the correlated random graph models exhibits sharp algorithmic phase transitions, as we elaborate below:
@@
  • For the detection problem between a pair of correlated Erdős-Rényi models $\mathcal{G}(n,q;\rho)$ and two independent Erdős-Rényi models $\mathcal{G}(n,q)$, we focus on the sparse regime where $q=n^{-1+o(1)}$. In this regime, on the one hand, it was shown in [MWXY24] that when $\rho>\sqrt{\alpha}$ where $\alpha\approx 0.338$ is the Otter’s constant [Ott48], there is an efficient algorithm that strongly distinguish these two models; on the other hand, it was shown in [DDL23+] that when $s<\sqrt{\alpha}$ there are evidences suggesting that all algorithms based on low-degree polynomials fail to strongly distinguish these two models.
@@
  • For the detection problem between a pair of correlated SBMs $\mathcal{S}(n,\frac{\lambda}{n};k,\epsilon;s)$ and two independent Erdős-Rényi graphs $\mathcal{G}(n,\frac{\lambda s}{n})$, we focus on the constant degree regime where $\lambda=O(1)$. In this regime, on the one hand, it was shown in [CDGL24+] that when $s>\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ where $\alpha\approx 0.338$ is the Otter’s constant and $\tfrac{1}{\epsilon^{2}\lambda}$ is the Kesten-Stigum threshold [KS66], there is an efficient algorithm that strongly distinguish these two models; on the other hand, it was also shown in [CDGL24+] that when $s<\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ there are evidences suggesting that all algorithms based on low-degree polynomials fail to strongly distinguish these two models.
@@
The lower bound in the aforementioned results explored inherent computational barriers from the perspective of the low-degree polynomial framework. Indeed, it has been proved that the class of low-degree polynomial algorithms is a useful proxy for computationally efficient algorithms, in the sense that the best-known polynomial-time algorithms for a wide variety of high-dimensional inference problems are captured by the low-degree class; see e.g. [HKP+17, Hop18, SW22, KWB22]. However, these aforementioned results suffer from two significant limitations, which we now discuss. Firstly, the aforementioned result cannot provide the evidence that partial matching recovery (that is, recover a positive fraction of the coordinates of the latent matching $\pi_{*}$) is impossible by efficient algorithms although efficient detection has already been ruled out. Secondly, (in the case of correlated SBM) they are only able to establish the computation threshold on the detection problem between the correlated model and a pair of independent Erdős-Rényi graphs. These limitations motivate two natural questions:
@@
    Can we provide the evidences that partial matching recovery is impossible (by efficient algorithms) in the same parameter regime where there are evidences suggesting that detection is impossible (by efficient algorithms)?
@@
The aim of this paper is to (partially) answer these two problems. Our main result can be informally summarized as follows:
@@
We have the following results hold.
@@
    Assuming a natural strengthening of the low-degree conjecture (see Conjecture 2.5 for its precise meaning), then for the correlated Erdős-Rényi model $\mathcal{G}(n,q,\rho)$, when $q=n^{-1+o(1)}$ and $\rho<\sqrt{\alpha}$ it is impossible to recover a positive fraction of the coordinates of $\pi_{*}$ with constant probability by efficient algorithms.
@@
    Assuming a natural strengthening of the low-degree conjecture (see Conjecture 2.5 for its precise meaning), then for the correlated stochastic block models $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$, when $\lambda=O(1)$ and $s<\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ it is impossible to recover a positive fraction of the coordinates of $\pi_{*}$ with probability tending to $1$ as $n\to\infty$ by efficient algorithms.
@@
    Assuming the low-degree conjecture (see Conjecture 2.2 for its precise meaning), then for the correlated stochastic block models $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon;s)$, when $s<\min\{\sqrt{\alpha},\tfrac{1}{\epsilon^{2}\lambda}\}$ it is impossible to strongly distinguish this model and a pair of independent SBMs $\mathcal{S}(n,\tfrac{\lambda s}{n};k,\epsilon)$ by efficient algorithms, provided that the average degree $\lambda s$ is sufficiently large.
@@
Note that when $\rho>\sqrt{\alpha}$, the results in [MWXY23, GM20, GML24, GMS24] show that there exists an efficient algorithm that achieves partial recovery of $\pi_{*}$ in a pair of correlated Erdős-Rényi models with probability tending to $1$ as $n\to\infty$. Thus, Item (1) in Theorem 1.6 is tight in some sense and the algorithmic partial recovery threshold is indeed given by $\rho=\sqrt{\alpha}$.
Similarly, the results in [MWXY23, GM20, GML24, GMS24] naturally extends to show that there exists an efficient algorithm that achieves partial recovery of $\pi_{*}$ in a pair of correlated SBMs when $s>\sqrt{\alpha}$ with probability tending to $1$ as $n\to\infty$. Thus, Item (2) in Theorem 1.6 is also tight in the subcritical regime (i.e., when marginally two graphs are below the KS threshold).
@@
From Item (3) in Theorem 1.6 we see that for a pair of correlated SBMs $(A,B)$, when marginally both $A$ and $B$ are below the KS threshold, there are evidences suggesting that no efficient algorithm can strongly distinguish $(A,B)$ from a pair of independent stochastic block models when the correlation $s<\sqrt{\alpha}$. Since the result in [MWXY24] extends naturally to the case of stochastic block models which provides an efficient algorithm that strongly distinguish these two models when $s>\sqrt{\alpha}$, we see that in the subcritical regime (i.e., when marginally both $A$ and $B$ are below the KS threshold) the algorithmic correlation detection threshold is given by $\sqrt{\alpha}$. On the contrary, in the supercritical regime where $A$ and $B$ are above the KS threshold, we believe that the algorithmic correlation detection threshold should be strict lower than $\sqrt{\alpha}$.
@@ 1.1 Key challenges and innovations
In this subsection, we briefly discuss our approach of proving Theorem 1.6 and some conceptual innovations behind it. Our idea can be informally summarized as follows:
@@
    We will show that the low-degree advantage between $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ and $\mathbb{Q}$ is bounded by an absolute constant. Then, from the standard low-degree conjecture (see Conjecture 2.2 for details) these two measures cannot be strongly distinguished by efficient algorithms. Thus, since $g_{i}$ is “of positive constant order” under the measure $\mathbb{P}(\cdot\mid\pi_{*}(i)=1)$ (as it should approximate $\mathbf{1}_{\{\pi_{*}(1)=i\}}$ in some sense) we expect that $g_{i}$ should also be “of positive constant order” under $\mathbb{Q}$.
@@
    Define $g=g_{1}+\ldots+g_{n}$. We then have a statistics $g$ that can be efficiently computed, and (I) under $\mathbb{Q}$ we expect that $g$ is “large” as it is the sum of $n$ “of positive constant order” terms; (II) under $\mathbb{P}$ we expect that $g$ is “not large” as each $g_{i}$ should approximate $\mathbf{1}_{\{\pi_{*}(1)=i\}}$ in some sense. Thus, the statistics $g$ accumulates more signals than all low-degree polynomials, which (non-rigorously speaking) violates the low-degree conjecture.
@@
    As for Item (3), denote $\mathbb{P}$ to be the law of correlated SBMs and $\mathbb{Q}$ of independent SBMs. Also denote $\widetilde{\mathbb{Q}}$ to be the law of independent Erdős-Rényi graphs. As it was already shown in [CDGL24+] that $\mathbb{P}$ and $\widetilde{\mathbb{Q}}$ cannot be strongly distinguished, (non-rigorously speaking) we only need to show that $\mathbb{Q}$ and $\widetilde{\mathbb{Q}}$ are also “indistinguishable” in some sense.
However, there are certain obstacles when implementing the above ideas, as we shall discuss below. In Step (2) we need to “transfer” the behavior of a statistics $g_{i}$ under $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ to its behavior under $\mathbb{Q}$. If the low-degree advantage between $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ and $\mathbb{Q}$ is bounded by $1+o(1)$, then the standard low-degree conjecture implies that we cannot distinguish between $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ and $\mathbb{Q}$ better than random by efficient algorithms. Consequently, we expect the behavior of $g_{i}$ to be “almost identical” under both $\mathbb{P}(\cdot\mid\pi_{*}(1)=i)$ and $\mathbb{Q}$ (as $g_{i}$ can be efficiently computed). However, in our case the low-degree advantage is just bounded by a large (but fixed) constant. This weaker condition means that the standard low-degree conjecture only rules out the possibility to distinguish these two measures efficiently with vanishing errors, leaving room for non-negligible distinctions. Similar issues also arise in Step (4), in which simply showing the low-degree advantage between $\mathbb{P}$ and $\widetilde{\mathbb{Q}}$ and the low-degree advantage between $\mathbb{Q}$ and $\widetilde{\mathbb{Q}}$ is bounded is not strong enough for our goal.
@@
Another difficulty is that reducing between different inference tasks requires a framework that better constrains the behavior of efficiently computable estimators. To address this, we introduce a natural refinement of the low-degree conjecture: we posit that low-degree polynomials achieve signal-to-noise ratios that are at least as good as those of any efficient algorithm (see Conjecture 2.5 for a more precise description). While similar modifications to the low-degree conjecture have been explored in [MW23+], our motivations differ significantly. Their work focused on characterizing the limits of precise error rate for all computationally feasible algorithms, whereas our goal is to enable reductions between distinct inference tasks.
1.2 Comparison to concurrent work
In a concurrent work [DHSS25+], the authors investigate the computational hardness of the weak recovery problem in the stochastic block model $\mathcal{S}(n,\tfrac{\lambda}{n};k,\epsilon)$ when $k=n^{o(1)}$ and $\epsilon^{2}\lambda<1$ lies below the KS threshold. Notably, their proof is also based on a “recovery-to-detection reduction” approach, which argues that if there exists an efficient algorithm that achieves weak recovery, then we can use this algorithm to efficiently construct a statistics for detection (in a certain sense). Their results also relies on a strengthening of the low-degree conjecture, which is similar to the assumptions made in this paper.
However, the reduction technique employed in this paper differs significantly from that of [DHSS25+]. The authors of [DHSS25+] construct their statistic for detection using a correlation preserving projection technique established in [HS17]. This technique, which is designed specifically for weak recovery in block models, enables them to “regularize” the statistics and thus bound its moments under the null hypothesis. In contrast, our construction is more straightforward as we leverage the universal framework described in Subsection 1.1 to “transfer” the behavior of statistics across different probability measures. It seems that in comparison to [DHSS25+], our approach is easier to implement, less dependent on problem-specific details, and potentially applicable to a broader range of problems.
1.3 Notations